# Investigation of ArUco Marker Placement for Planar Indoor Localization

### v2

@@ Abstract
Note: This is the accepted version of the paper. The final version will be published in the Proceedings of the IEEE 2025 22nd International Conference on Advanced Robotics (ICAR).
@@ III-A Pinhole camera model
For camera pose estimation, this work makes use of the pinhole camera model [31, p. 153ff.], which assumes that an image is generated by rays of light passing through the camera’s pinhole and hitting the sensor plane consisting of photo diodes. We define a 3D point in world coordinates as $\underline{p}_{w}=[x_{w},y_{w},z_{w}]^{T}\in\mathbb{R}^{3}$. The corresponding homogeneous coordinates are $\underline{\overline{p}}_{w}=[x_{w},y_{w},z_{w},1]^{T}\in\mathbb{R}^{4}$. Its 2D projection in image coordinates $\underline{p}_{i}=[x_{i}/z_{i},y_{i}/z_{i}]^{T}\in\mathbb{R}^{2}$ can be computed from the homogeneous representation $\overline{\underline{p}}_{i}=[x_{i},y_{i},z_{i}]^{T}\in\mathbb{R}^{3}$. The projection of corresponding points in world coordinates $\underline{\overline{p}}_{w}$ to image coordinates $\underline{\overline{p}}_{i}$ is given by
@@ III-B Perspective-n-Point problem
For camera-based localization, the PnP problem is the problem of estimating extrinsic parameters from known correspondences between $n$ 2D image points and $n$ 3D world points [31, p. 178f.]. Generally, the goal of the estimation is to find the 12 entries of the projection matrix $\mathbf{P}$ [31, p. 156] to determine the 6-DoF camera pose (3 parameters for the translation vector and 3 independent rotations from the 3x3 rotation matrix) such that the reprojection error based on the Euclidean distances between the detected 2D image points $\underline{\hat{p}}_{ik}$ and the 2D image projections $\underline{p}_{ik}$ from corresponding 3D world points is minimized [31, p. 181]. From this, the 6 extrinsic parameters can be extracted if the intrinsic matrix $\mathbf{K}$ is known. For a general solution of the PnP problem, at least $n=6$ pairs $(\underline{p}_{ik},\underline{\hat{p}}_{ik})$ are needed, though there exist solvers that are able to work with less points, if they fulfill additional criteria, allowing for 6-DoF camera pose estimation with a single marker. The solvers supported by OpenCV as well as their requirements can be found in the OpenCV documentation for SolvePnPMethod [10]. For this work, the ITERATIVE scheme is used in combination with random sample consensus (RANSAC) to gain more resistance to outliers. The solver itself is based on direct linear transform (DLT) to find an initial estimate for $\mathbf{P}$ and a subsequent Levenberg-Marquardt optimization.
@@ III-C Camera calibration
The intrinsic matrix $\mathbf{K}$ is determined offline through camera calibration. During localization, $\mathbf{K}$ is passed to the OpenCV PnP solver to obtain the extrinsic parameters. Camera calibration requires at least ten images from different perspectives of a known calibration pattern. The images are then used to first estimate the homography for each view and then the intrinsic parameters [32, p. 665f.]. Commonly used calibration patterns include a chessboard, ArUco or ChArUco patterns [32, p. 652ff.]. During this work, the chessboard pattern was used, as it led to the best pose estimates. Calibration with a liquid-crystal display (LCD) instead of a printed version, as suggested in [33], was examined and did not bring improvements in localization.
@@ III-E Random Sampling Consensus (RANSAC)
Given a set of data points, the RANSAC algorithm randomly selects a minimum number of them to solve for the (extrinsic) parameters. In a second step, the remaining set is divided into inliers and outliers, based on whether they fit the estimated model or not [31, p. 117ff.]. The OpenCV solvePnPRansac function [10] makes use of this by randomly selecting a minimum number of correspondences, calling a PnP solver for the selected pairs, and computing the reprojection errors to divide all samples into inliers and outliers. After repeating the procedure of random sampling, model fitting, and inlier/outlier division for a given number of iterations, RANSAC returns the model with the most inliers. RANSAC can deal with a large number of outliers, thus improving the robustness of localization.
@@ III-F Camera pose estimation
Using the detected marker corners and their measured positions in the world coordinate system as well as the intrinsic parameters gained by camera calibration, the solver can be called to estimate $\mathbf{P}$ and directly extract $\mathbf{R}$ and $\underline{t}$. Once the extrinsic parameters are extracted from $\mathbf{P}$, the 6-DoF camera pose (i.e. it’s position $\underline{C}\in\mathbb{R}^{3}$ and the 3 Euler angles including the heading $\theta$, derived from the rotation matrix $\mathbf{R}_{C}\in\mathbb{R}^{3\times 3}$) is computed with [31, p. 163f.]
@@ IV-A Kalman filter for fiducial marker based localization
In this work, $\mathbf{Q}_{x}\in\mathbb{R}^{2\times 2}$ and $\mathbf{Q}_{y}\in\mathbb{R}^{2\times 2}$ are defined using a discrete white noise acceleration (DWNA) model, which assumes that the constant velocity system is disturbed by a piecewise constant acceleration, where the accelerations per time step are uncorrelated [35, p. 273]. The variances for the accelerations in $x$ and $y$ are $\sigma^{2}_{a_{x}}$ and $\sigma^{2}_{a_{y}}$, respectively. We experimentally set $\sigma^{2}_{a_{x}}=\sigma^{2}_{a_{y}}=1\left(\frac{\mathrm{cm}}{\mathrm{s}^{2}}\right)^{2}$ such that measurements are trusted more than the motion model.

### v3

@@ Abstract
Note: This is the accepted version of the paper. The final version is published in the Proceedings of the IEEE 2025 22nd International Conference on Advanced Robotics (ICAR). DOI: 10.1109/ICAR65334.2025.11338671
@@ III-A Pinhole camera model
For camera pose estimation, this work makes use of the pinhole camera model [12, p. 153ff.], which assumes that an image is generated by rays of light passing through the camera’s pinhole and hitting the sensor plane consisting of photo diodes. We define a 3D point in world coordinates as $\underline{p}_{w}=[x_{w},y_{w},z_{w}]^{T}\in\mathbb{R}^{3}$. The corresponding homogeneous coordinates are $\underline{\overline{p}}_{w}=[x_{w},y_{w},z_{w},1]^{T}\in\mathbb{R}^{4}$. Its 2D projection in image coordinates $\underline{p}_{i}=[x_{i}/z_{i},y_{i}/z_{i}]^{T}\in\mathbb{R}^{2}$ can be computed from the homogeneous representation $\overline{\underline{p}}_{i}=[x_{i},y_{i},z_{i}]^{T}\in\mathbb{R}^{3}$. The projection of corresponding points in world coordinates $\underline{\overline{p}}_{w}$ to image coordinates $\underline{\overline{p}}_{i}$ is given by
@@ III-B Perspective-n-Point problem
For camera-based localization, the PnP problem is the problem of estimating extrinsic parameters from known correspondences between $n$ 2D image points and $n$ 3D world points [12, p. 178f.]. Generally, the goal of the estimation is to find the 12 entries of the projection matrix $\mathbf{P}$ [12, p. 156] to determine the 6-DoF camera pose (3 parameters for the translation vector and 3 independent rotations from the 3x3 rotation matrix) such that the reprojection error based on the Euclidean distances between the detected 2D image points $\underline{\hat{p}}_{ik}$ and the 2D image projections $\underline{p}_{ik}$ from corresponding 3D world points is minimized [12, p. 181]. From this, the 6 extrinsic parameters can be extracted if the intrinsic matrix $\mathbf{K}$ is known. For a general solution of the PnP problem, at least $n=6$ pairs $(\underline{p}_{ik},\underline{\hat{p}}_{ik})$ are needed, though there exist solvers that are able to work with less points, if they fulfill additional criteria, allowing for 6-DoF camera pose estimation with a single marker. The solvers supported by OpenCV as well as their requirements can be found in the OpenCV documentation for SolvePnPMethod [8]. For this work, the ITERATIVE scheme is used in combination with random sample consensus (RANSAC) to gain more resistance to outliers. The solver itself is based on direct linear transform (DLT) to find an initial estimate for $\mathbf{P}$ and a subsequent Levenberg-Marquardt optimization.
@@ III-C Camera calibration
The intrinsic matrix $\mathbf{K}$ is determined offline through camera calibration. During localization, $\mathbf{K}$ is passed to the OpenCV PnP solver to obtain the extrinsic parameters. Camera calibration requires at least ten images from different perspectives of a known calibration pattern. The images are then used to first estimate the homography for each view and then the intrinsic parameters [15, p. 665f.]. Commonly used calibration patterns include a chessboard, ArUco or ChArUco patterns [15, p. 652ff.]. During this work, the chessboard pattern was used, as it led to the best pose estimates. Calibration with a liquid-crystal display (LCD) instead of a printed version, as suggested in [29], was examined and did not bring improvements in localization.
@@ III-E Random Sampling Consensus (RANSAC)
Given a set of data points, the RANSAC algorithm randomly selects a minimum number of them to solve for the (extrinsic) parameters. In a second step, the remaining set is divided into inliers and outliers, based on whether they fit the estimated model or not [12, p. 117ff.]. The OpenCV solvePnPRansac function [8] makes use of this by randomly selecting a minimum number of correspondences, calling a PnP solver for the selected pairs, and computing the reprojection errors to divide all samples into inliers and outliers. After repeating the procedure of random sampling, model fitting, and inlier/outlier division for a given number of iterations, RANSAC returns the model with the most inliers. RANSAC can deal with a large number of outliers, thus improving the robustness of localization.
@@ III-F Camera pose estimation
Using the detected marker corners and their measured positions in the world coordinate system as well as the intrinsic parameters gained by camera calibration, the solver can be called to estimate $\mathbf{P}$ and directly extract $\mathbf{R}$ and $\underline{t}$. Once the extrinsic parameters are extracted from $\mathbf{P}$, the 6-DoF camera pose (i.e. it’s position $\underline{C}\in\mathbb{R}^{3}$ and the 3 Euler angles including the heading $\theta$, derived from the rotation matrix $\mathbf{R}_{C}\in\mathbb{R}^{3\times 3}$) is computed with [12, p. 163f.]
@@ IV-A Kalman filter for fiducial marker based localization
In this work, $\mathbf{Q}_{x}\in\mathbb{R}^{2\times 2}$ and $\mathbf{Q}_{y}\in\mathbb{R}^{2\times 2}$ are defined using a discrete white noise acceleration (DWNA) model, which assumes that the constant velocity system is disturbed by a piecewise constant acceleration, where the accelerations per time step are uncorrelated [4, p. 273]. The variances for the accelerations in $x$ and $y$ are $\sigma^{2}_{a_{x}}$ and $\sigma^{2}_{a_{y}}$, respectively. We experimentally set $\sigma^{2}_{a_{x}}=\sigma^{2}_{a_{y}}=1\left(\frac{\mathrm{cm}}{\mathrm{s}^{2}}\right)^{2}$ such that measurements are trusted more than the motion model.